a rectangular floor 12 units long and 8 units wide is to be covered w/ triangular tiles w/ side 3 units, 4 units and 5 units how many tiles are neede?

Answer by god2012(113)

the sides of the triangle are 3,4 and 5 units
That is , it is a right angled triangle (pythagorian triplets)
Area of triangle = 1/2 * 4 * 3 = 6 sq. units
Area of rectangle = L * B = 12 * 8 = 96 Sq. units
Total number of tiles required = total area / total area of triangle per tile
= 96/6 = 16 tiles
Total tiles needed to cover the floor = 16